Two samples of n = 5 subjects were analyzed with an independent-samples t-test, producing a t = 2.295.

         a) What is the tcrit ?

         b) Does this indicate a significant difference in a two-tailed test with α = .05?
Answer #1

Part a

We are given

Test is two tailed test.

n1 = 5

n2 = 5

df = n1 + n2 – 2 = 5 + 5 – 2 = 8

df = 8

α = 0.05

So, critical value by using t-table or excel is given as below:

tcritcial = ±2.3060

Part b

We have

tcritcial = ±2.3060

tobs= 2.295

tobs = 2.295 < tcritcial = 2.3060

So, we do not reject the null hypothesis

This does not indicate a significant difference in a two-tailed test with α = 0.05.